After five half lives, percentage of original radioactive atoms left is
1%
0.3%
3.125%
0.2%
 

Grade:12

3 Answers

Sumit Majumdar IIT Delhi
8 years ago
We have:
N=N_0\exp^{-\lambda t}
And,
\lambda=\frac{0.693}{t_{1/2}}
Hence,
N=N_0\exp^{-\lambda t}=N_0\exp^{-\frac{0.693}{t_{1/2}} \times {5t_{1/2}}}=N_0\exp^{-{0.693}\times {5}}=N_0\exp^{-3.465}
So,
\frac{N}{N_0}=3.127 \%

Adarsh
8 years ago
After 1, 2, 3, 4 half lives the percentage of original radioactive atoms left will be
50%, 25%, 12.5%, 6.25%.
So after 5 half lives the percentage of original radioactive atoms left will be
3.125%
Please approve if my answer is correct
prince Kumar
6 years ago
6 will be answer after 4 half lifes of radioactive atom this is right or wrong tell me. .who discovered heavy hydrogen?